A simple pendulum makes 5 oscillations in 1 minute determine the oscillation
1
Expert's answer
2019-09-27T10:00:39-0400

Since the words "Determine the oscillations" have uncertain meaning, determine the period, frequency and the angular frequency.

A period is time required to make 1 complete oscillation (convert time to seconds!):


"T=\frac{t}{N}=\frac{60}{5}=12\text{ s}."

A frequency is how many oscillations the pendulum makes in a second:


"f=\frac{1}{T}=\frac{1}{12}=0.083\text{ Hz}."

An angular frequency is an important value which allows determining different mathematical characteristics of oscillations:


"\omega=2\pi f=2\cdot3.14\cdot0.083=0.52\text{ s}^{-1}."